OpenStudy (anonymous):

Can any of you guys help me !?? Simplify square root of negative 72. a. negative 6 square root of 2 b. 6 square root of negative 2 c. 6 square root of 2 i d. 6 I square root of 2

OpenStudy (anonymous):

first do you know what is the \[\sqrt{-1}\]

OpenStudy (anonymous):

yeah .. i dnt want u to give me the answer i just need help on how to solve them.

OpenStudy (anonymous):

what are the factors in 72? Can you break it down?

OpenStudy (anonymous):

its D right?

OpenStudy (anonymous):

I presume D is \[6i \sqrt{2}\]

OpenStudy (anonymous):

yeah..

OpenStudy (anonymous):

In order to solve this you need to understand the concept of i. In math, there is an imaginary number called "i" and it usually comes up when you are simplifying negative radicals.

OpenStudy (anonymous):

that you got it correct

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

Create an equation that results in at least one extraneous solution. Work through your equation, justify each step, and explain how the solution is extraneous. Using the properties of exponents and radicals, design at least three different equivalent forms of x². You must show how each one can be simplified back to x² in two or more steps. Stretch your mind and get creative! Keep in mind that something too simple, like x • x would not be acceptable since it takes only one step to convert it to x². Substitute your birthday into the equation Square root of the quantity x minus y + m = d, where y is the last two digits of your birth year, m is the month, and d is the day. If you were born on 7/10/1856 like Nikola Tesla, your equation would be Square root of the quantity x minus 56 + 7 = 10. Solve for x and identify if it is an extraneous solution.

OpenStudy (anonymous):

@WordGEEK can you help?

OpenStudy (anonymous):

Please put it as a new question.